# Trust-Aware Assistance-Seeking in Human-Supervised Autonomy

### v2

@@ 4.3. Model Parameter Confidence
We now investigate uncertainty in the estimated parameters. Since the Baum-Welch algorithm provides a maximum a posteriori estimate, we rely on Laplace approximation (pml2Book, ) to estimate parameter uncertainty. To this end, we estimate the Hessian matrix $H$ of the log-likelihood $L$ and compute the uncertainty in parameter estimates as $e=\sqrt{\operatorname{diag}(-H)}$, where $\operatorname{diag}$ represents the diagonal elements of the matrix H-H. The uncertainties in the estimated IOHMM human trust model parameters are shown in Table 1.
@@ 4.4. Trust as a Hidden State
We now investigate if it is justified to refer to the hidden state in our model as “human trust”. During the experiment, after each subblock, participants completed a survey (lee1992trust, ) in which they responded, on an 11-point scale, to the following question “To what extent do you trust the automation in this scenario?”, shown in Appendix 8.2. This question was intended to capture participants’ level of trust based on their ongoing experience while performing the task. This survey is collected each time the robot finishes collecting objects from a shelf, before moving to another shelf.
@@ 5.3. Survey-Based Evaluation of the Trust-Aware Policy
We now investigate if the trust-aware policy improved human trust in addition to improving the team performance as discussed above. In our experiments, participants completed the Trust Perception Scale survey both before and after the experiment. The Trust Perception Scale is a 40-item survey that aims to measure trust while taking into account the human, robot, and environmental elements of the human-robot interaction. It is designed as a pre- and post-interaction measurement of trust and is used to evaluate changes in human trust (schaefer2016measuring, ). The collected responses are the 14 sub-scale of the survey developed in (schaefer2016measuring, ), also shown in Appendix 8.1.
In total, participants completed this questionnaire three times, the first before the experiment, and the remaining two after completion with each policy. We dub these survey responses as “pre-experiment”, “post-trust-agnostic”, and “post-trust-aware” responses. The comparison of the survey scores is shown in Fig. 12. The responses to the 14 survey questions are aggregated, resulting in a trust score. This process involves first reverse coding the questions that describe negative qualities. The responses are then summed and divided by the total number of questions. More detailed instructions can be found in (schaefer2016measuring, ). The average scores on the collected trust survey are 65.44, 66.39, and 72.14 for pre-experiment, trust-agnostic, and trust-aware, respectively.
@@ 6. Discussion
The concept of trust has been explored across various domains, including both human-human and human-robot interactions. Mayer et al. (mayer1995integrative, ) formally define trust between individuals as “the willingness of a party to be vulnerable to the outcomes of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party.”.
@@
The definition emphasizes that trust becomes crucial when vulnerability, risk, or uncertainty is involved in an interaction. With the deployment of robots working in close proximity with human partners in risky and uncertain environments (takayama2008beyond, ), human trust in robots becomes a crucial factor for their adoption. While human-robot trust shares commonality with human-human trust, the former is typically considered only from the human’s perspective regarding the system’s performance (baker2018toward, ). However, researchers have begun exploring bidirectional trust in human-robot interactions (rahman2018mutual, ; Wang2014, ).
@@
We developed an IOHMM model for trust-modulated human behavior and estimated its parameters using data collected from the experiment. Along with the definition of trust in automation, Lee and See (lee2004trust, ), provided a conceptual model of trust, describing the process and interaction of factors when taking trust into account. Our IOHMM model complies with Lee and See’s description of how factors interact in the dynamic process when taking trust into account (lee2004trust, ). In our model, the information assimilation is represented by the input to the IOHMM, and the trust evolution is governed by the state transition probabilities. Intention formation is decided by the emission probabilities and human action is represented by the IOHMM output. Our fitted model suggests that success can increase trust, while failure can decrease it. The reliability of the system is a key factor that shapes human trust in robots (hancock2011meta, ). It is well known that positive experiences or successes increase human trust, while negative experiences or failures decrease it (chen2020trust, ; akash2020human, ; azevedo2021real, ). Additionally, humans tend to rely more on robots when less risk is involved (chen2020trust, ; akash2020human, ). As expected, our model indicates that success in high-complexity tasks has a greater impact on increasing trust compared to success in low-complexity tasks.
@@
We investigated the robot’s action of requesting assistance from human supervisors, enabling the robot to transfer control of the current task to the human supervisor. Our model estimates suggest that asking for assistance in high-complexity tasks can help increase trust. It has been shown that humans are more trusting of risk-averse robots (bridgwater2020examining, ). When human trust is low, their belief that the robot will fail is high, and by seeking assistance, humans perceive the robot as being cautious and avoiding potential failure by not attempting to collect autonomously.
@@
We grounded our model by confirming that its hidden state is isomorphic to self-reported human trust. Trust model parameters have been estimated using self-reported trust from human participants (chen2020trust, ; akash2017dynamic, ; zahedi2023trust, ; mangalin2024, ). Besides reported trust levels, models have also been estimated using behavioral data, such as action and task performance, from human participants (azevedo2021real, ; akash2020human, ; DBLP:journals/corr/abs-2009-11890, ; mangalin2025trustrep, ; williams2023computational, ). Traditionally, when human behavior data is used for model parameter estimation, the state of the system is labeled as the human trust but most work does not directly justify this. In our work, we have verified that our trust model’s estimate is isomorphic to participants’ trust survey responses. With this, we demonstrated that our model accurately captures the dynamics of human trust based solely on behavioral data, without relying on survey responses.
@@
To investigate human trust accurately, a suitable trust-elicitation experiment must be designed to capture trust dynamics effectively. The most prominent experimental paradigm to study human-robot trust relies on block arrangements, wherein the robot is highly reliable in one block and less reliable in the other (salem2015would, ; manzey2012human, ). While these paradigms provide insights into human-robot trust, they do not generate rich enough data to build dynamic models of trust. To address this challenge, we adopted a randomized experimental design wherein the task complexity, influencing robot success rate, is randomly selected in each trial. This approach more naturally excites various levels of human trust and its interaction with different task outcomes, including unexpected and random failures that humans experience when working with robots. Additionally, we believe this experimental paradigm is a closer representation of real-world human-robot interaction scenarios, and is being increasingly adopted in HRI studies (akash2020human, ; azevedo2021real, ; DBLP:journals/corr/abs-2009-11890, )
@@
The IOHMM/POMDP-based framework provides a formal structure for inferring human cognitive state and designing optimal policies for stochastic human-robot systems and is being increasingly used in the HRI research community. Despite their advantages, the data requirement to fit an IOHMM, and the computational requirements to compute an optimal policy for the POMDP often limit the size of the state space, for example, most trust models leveraging this framework use a binary trust state (wang2023human, ). A potential approach to address these issues is to rely on difference equation models (jonker1999formal, ) and rely on techniques such as Model Predictive Control (rawlings2017model, ) to design the optimal policy (mangalin2024, ).
@@
In this work, the computed optimal assistance-seeking policy is a threshold-based policy, robot seeks assistance only when its belief that the human trust is low is above a threshold. Sufficient conditions have been established for an optimal policy for a POMDP to admit such a structure (krishnamurthy2016partially, ; gupta2024structural, ). Future research will leverage these results to generic operational characteristics of the robot that result in such policies.

### v3

@@ 4.3. Model Parameter Confidence
We now investigate uncertainty in the estimated parameters. Since the Baum-Welch algorithm provides a maximum a posteriori estimate, we rely on Laplace approximation (pml2Book) to estimate parameter uncertainty. To this end, we estimate the Hessian matrix $H$ of the log-likelihood $L$ and compute the uncertainty in parameter estimates as $e=\sqrt{\operatorname{diag}(-H)}$, where $\operatorname{diag}$ represents the diagonal elements of the matrix H-H. The uncertainties in the estimated IOHMM human trust model parameters are shown in Table 1.
@@ 4.4. Trust as a Hidden State
We now investigate if it is justified to refer to the hidden state in our model as “human trust”. During the experiment, after each subblock, participants completed a survey (lee1992trust) in which they responded, on an 11-point scale, to the following question “To what extent do you trust the automation in this scenario?”, shown in Appendix 8.2. This question was intended to capture participants’ level of trust based on their ongoing experience while performing the task. This survey is collected each time the robot finishes collecting objects from a shelf, before moving to another shelf.
@@ 5.3. Survey-Based Evaluation of the Trust-Aware Policy
We now investigate if the trust-aware policy improved human trust in addition to improving the team performance as discussed above. In our experiments, participants completed the Trust Perception Scale survey both before and after the experiment. The Trust Perception Scale is a 40-item survey that aims to measure trust while taking into account the human, robot, and environmental elements of the human-robot interaction. It is designed as a pre- and post-interaction measurement of trust and is used to evaluate changes in human trust (schaefer2016measuring). The collected responses are the 14 sub-scale of the survey developed in (schaefer2016measuring), also shown in Appendix 8.1.
In total, participants completed this questionnaire three times, the first before the experiment, and the remaining two after completion with each policy. We dub these survey responses as “pre-experiment”, “post-trust-agnostic”, and “post-trust-aware” responses. The comparison of the survey scores is shown in Fig. 12. The responses to the 14 survey questions are aggregated, resulting in a trust score. This process involves first reverse coding the questions that describe negative qualities. The responses are then summed and divided by the total number of questions. More detailed instructions can be found in (schaefer2016measuring). The average scores on the collected trust survey are 65.44, 66.39, and 72.14 for pre-experiment, trust-agnostic, and trust-aware, respectively.
@@ 6. Discussion
The concept of trust has been explored across various domains, including both human-human and human-robot interactions. Mayer et al. (mayer1995integrative) formally define trust between individuals as “the willingness of a party to be vulnerable to the outcomes of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party.”.
@@
The definition emphasizes that trust becomes crucial when vulnerability, risk, or uncertainty is involved in an interaction. With the deployment of robots working in close proximity with human partners in risky and uncertain environments (takayama2008beyond), human trust in robots becomes a crucial factor for their adoption. While human-robot trust shares commonality with human-human trust, the former is typically considered only from the human’s perspective regarding the system’s performance (baker2018toward). However, researchers have begun exploring bidirectional trust in human-robot interactions (rahman2018mutual; Wang2014).
@@
We developed an IOHMM model for trust-modulated human behavior and estimated its parameters using data collected from the experiment. Along with the definition of trust in automation, Lee and See (lee2004trust), provided a conceptual model of trust, describing the process and interaction of factors when taking trust into account. Our IOHMM model complies with Lee and See’s description of how factors interact in the dynamic process when taking trust into account (lee2004trust). In our model, the information assimilation is represented by the input to the IOHMM, and the trust evolution is governed by the state transition probabilities. Intention formation is decided by the emission probabilities and human action is represented by the IOHMM output. Our fitted model suggests that success can increase trust, while failure can decrease it. The reliability of the system is a key factor that shapes human trust in robots (hancock2011meta). It is well known that positive experiences or successes increase human trust, while negative experiences or failures decrease it (chen2020trust; akash2020human; azevedo2021real). Additionally, humans tend to rely more on robots when less risk is involved (chen2020trust; akash2020human). As expected, our model indicates that success in high-complexity tasks has a greater impact on increasing trust compared to success in low-complexity tasks.
@@
We investigated the robot’s action of requesting assistance from human supervisors, enabling the robot to transfer control of the current task to the human supervisor. Our model estimates suggest that asking for assistance in high-complexity tasks can help increase trust. It has been shown that humans are more trusting of risk-averse robots (bridgwater2020examining). When human trust is low, their belief that the robot will fail is high, and by seeking assistance, humans perceive the robot as being cautious and avoiding potential failure by not attempting to collect autonomously.
@@
We grounded our model by confirming that its hidden state is isomorphic to self-reported human trust. Trust model parameters have been estimated using self-reported trust from human participants (chen2020trust; akash2017dynamic; zahedi2023trust; mangalin2024). Besides reported trust levels, models have also been estimated using behavioral data, such as action and task performance, from human participants (azevedo2021real; akash2020human; DBLP:journals/corr/abs-2009-11890; mangalin2025trustrep; williams2023computational). Traditionally, when human behavior data is used for model parameter estimation, the state of the system is labeled as the human trust but most work does not directly justify this. In our work, we have verified that our trust model’s estimate is isomorphic to participants’ trust survey responses. With this, we demonstrated that our model accurately captures the dynamics of human trust based solely on behavioral data, without relying on survey responses.
@@
To investigate human trust accurately, a suitable trust-elicitation experiment must be designed to capture trust dynamics effectively. The most prominent experimental paradigm to study human-robot trust relies on block arrangements, wherein the robot is highly reliable in one block and less reliable in the other (salem2015would; manzey2012human). While these paradigms provide insights into human-robot trust, they do not generate rich enough data to build dynamic models of trust. To address this challenge, we adopted a randomized experimental design wherein the task complexity, influencing robot success rate, is randomly selected in each trial. This approach more naturally excites various levels of human trust and its interaction with different task outcomes, including unexpected and random failures that humans experience when working with robots. Additionally, we believe this experimental paradigm is a closer representation of real-world human-robot interaction scenarios, and is being increasingly adopted in HRI studies (akash2020human; azevedo2021real; DBLP:journals/corr/abs-2009-11890)
@@
The IOHMM/POMDP-based framework provides a formal structure for inferring human cognitive state and designing optimal policies for stochastic human-robot systems and is being increasingly used in the HRI research community. Despite their advantages, the data requirement to fit an IOHMM, and the computational requirements to compute an optimal policy for the POMDP often limit the size of the state space, for example, most trust models leveraging this framework use a binary trust state (wang2023human). A potential approach to address these issues is to rely on difference equation models (jonker1999formal) and rely on techniques such as Model Predictive Control (rawlings2017model) to design the optimal policy (mangalin2024).
@@
In this work, the computed optimal assistance-seeking policy is a threshold-based policy, robot seeks assistance only when its belief that the human trust is low is above a threshold. Sufficient conditions have been established for an optimal policy for a POMDP to admit such a structure (krishnamurthy2016partially; gupta2024structural). Future research will leverage these results to generic operational characteristics of the robot that result in such policies.
@@ Acknowledgements
8. Appendices
This appendix presents the survey questions that were collected, additional experimental results, and a summary of the survey data that complements the findings discussed in the main text. The Trust Perception Scale (schaefer2016measuring) is presented in Section 8.1 along with a summary of participant responses. The trust and reliance survey collected throughout the experiment is shown in Section 8.2, accompanied by a statistical summary. The statistical summary of surveys indicates that participants reported higher levels of trust when interacting with the trust-aware policy compared to the trust-agnostic policy. Additionally, we conducted experiments in which participants’ compensation depended on team performance. The results showing both task performance and human trust are presented in Section 8.3.
8.1. Trust Perception Scale
The Trust Perception Scale is a 40-item survey that aims to measure trust while taking into account the human, robot, and environmental elements of the human-robot interaction. It is designed as a pre- and post-interaction measurement of trust and is used to evaluate changes in human trust (schaefer2016measuring). We have collected the 14 sub-scales of the survey:
We collected survey responses at the beginning of the experiment and after participants interacted with each of the two policies. The summaries of participants’ responses are presented in Table 2, which provides descriptive statistics for all survey items across each experimental condition. The results show that the Trust-Aware policy generally received higher mean ratings for positive characteristics (e.g., dependability, reliability) and lower ratings for negative characteristics (e.g., malfunction) compared to the Trust-Agnostic policy. Overall, the Trust-Aware policy received higher ratings than the Trust-Agnostic policy.
Item Pre-Experiment Trust-Aware Trust-Agnostic
Mean Median Variance Mean Median Variance Mean Median Variance
1 64.44 70.00 226.14 76.11 80.00 154.58 69.44 70.00 217.32
2 70.00 70.00 152.94 75.00 80.00 179.41 69.44 70.00 182.03
3 33.89 30.00 507.52 21.11 10.00 739.87 18.33 10.00 520.59
4 67.78 70.00 430.07 77.22 80.00 127.12 70.56 70.00 229.08
5 69.44 70.00 264.38 81.67 85.00 120.59 72.22 80.00 218.30
6 63.89 70.00 225.16 78.89 80.00 139.87 71.67 70.00 238.24
7 37.22 25.00 844.77 22.78 10.00 562.42 27.22 25.00 515.36
8 33.89 25.00 578.10 20.56 10.00 440.85 28.89 25.00 469.28
9 43.33 50.00 1011.76 37.78 30.00 1041.83 32.22 20.00 1065.36
10 66.67 70.00 247.06 80.00 80.00 141.18 73.33 75.00 188.24
11 62.78 60.00 303.59 67.78 75.00 571.24 63.33 75.00 682.35
12 53.89 60.00 860.46 39.44 35.00 1029.08 28.33 20.00 967.65
13 71.11 70.00 316.34 76.67 80.00 200.00 73.33 80.00 223.53
14 76.67 80.00 400.00 83.89 90.00 154.58 80.00 80.00 329.41
Aggregate 64.64 65 138.79 72.14 72.50 68.19 66.39 69.29 54.17
Table 2. Combined descriptive statistics (mean, median, variance) for each survey item across Pre-Experiment, Trust-Agnostic, and Trust-Aware conditions. It can be seen that, when comparing the mean values for each item, in general, when looking at positive characteristics (i.e., dependability, reliability), the trust-aware policy received a higher mean rating compared to the trust-agnostic policy. Similarly, in items pertaining to negative characteristics (i.e., malfunction), the trust-aware policy received a lower rating compared to the trust-agnostic policy. Aggregating the results, the trust-aware policy received a higher rating compared to the trust-agnostic policy.
8.2. Trust and Reliance Survey
These surveys are collected after a participant finishes a subblock or shelf. They are used to measure self-reported human trust and reliance throughout the task. These survey questions are based on the seminal work (lee1992trust; lee1994trust).
A statistical summary of the survey responses collected is shown in Table 3, which compares participants’ ratings for the Trust-Aware and Trust-Agnostic policies. As shown, the Trust-Aware policy received higher trust ratings overall, while the reliance ratings for both policies were approximately the same. These results suggest that participants were generally more trusting of the Trust-Aware policy.
TRUST SURVEY Mean Median Variance
Trust-agnostic 6.68 7.33 4.19
Trust-aware 7.71 7.75 2.02
RELIANCE SURVEY Mean Median Variance
Trust-agnostic 7.66 7.83 2.92
Trust-aware 7.62 7.83 2.15
Table 3. Trust and reliance survey statistics for trust-aware and trust-agnostic policies. Surveys are collected during the policy evaluation experiment, and we show and compare the statistics for each policy. On average, the trust-aware policy received a higher rating in the trust survey compared to the trust-agnostic policy, while the reliance ratings for both policies were approximately the same. This further suggests that participants were more trusting of the trust-aware policy.
8.3. Supplemental experiment: Task performance and Trust Score
We conducted additional experiments with 6 participants, in which their compensation depended on team performance to further incite human trust. The average compensation received by the participants was $\mathdollar 13.18$. The results are shown in Fig. 13, and are consistent with the above results, i.e., the trust-aware policy outperformed the trust-agnostic policy, and the trust-agnostic policy received a higher trust score from participants.
(a) Cumulative reward statistics for the trust-agnostic and trust-aware policies. The trust-aware policy outperformed the trust-agnostic policy, consistent with the results discussed in Section 5.2.
(b) Comparison of the collected trust surveys before the experiment and after the experiment with the trust-aware and trust-agnostic policy. Results suggest that human trust is higher when working with trust-aware policy, consistent with the results discussed in Section 5.3.
Figure 13. Results from the supplemental experiment on (a) task performance and (b) trust score under performance-based compensation.: